OpenStudy (anonymous):

Alg 1 Honors What are the solutions of 6x2 = 12x? a) x = 0, x = −2 b) x = 0, x = 2 c) x = 6, x = 2 d) x = 6, x = −2 Thank You!

OpenStudy (anonymous):

6x2-12x=0 6x(x-2) x=0, x=2 b

OpenStudy (anonymous):

Thank You!:)

OpenStudy (anonymous):

No problem! :)

OpenStudy (anonymous):

Can you help me with a few more please? Identify the solutions of 3x2 + 14x + 8 = 0. x = −8 over 3, x = −1 x = 8 over 3, x = 1 x = −2 over 3, x = −4 x = 2 over 3, x = 4

OpenStudy (anonymous):

Factor (3x+2)(x+4) Solve C. X=-2/3, X= -4

OpenStudy (anonymous):

Thank You:))

OpenStudy (anonymous):

Anymore?